# Actively Learning Combinatorial Optimization Using a Membership Oracle

### v2

@@ 3.1 Closest Cutting Plane Sampling Strategy
For the pair of sets $S^{+}_{t}$ and $S^{-}_{t}$ of feasible and infeasible points at time step $t$, the current version space is defined as the set:
@@
Let us also define $S^{+}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1\}$ and $S^{-}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}>1\}$, as the sets of all feasible and infeasible points, respectively. Every time we update $S^{+}_{t}$ and $S^{-}_{t}$ with new labeled points to obtain $S^{+}_{t+1}$ and $S^{-}_{t+1}$, the next version space $V_{t+1}$ will be defined by additional inequalities with respect to $V_{t}$. The sequence of version spaces converges (in a finite but potentially exponential number of steps) to the set of surrogate weights whose corresponding surrogate constraint classifies every point in $X$ on the same side as the unknown knapsack constraint $\bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$. Namely,
@@
=$=\{\boldsymbol{w}\in W\ |\ \hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{+}\leq 1\ \text{for all}\ \boldsymbol{s}^{+}\in S^{+}_{\infty}\text{,}\ \boldsymbol{w}\cdot\boldsymbol{s}^{-}>1\ \text{for all}\ \boldsymbol{s}^{-}\in S^{-}_{\infty}\}\text{.}$ (3)
Intuitively, if the vector of surrogate constraints $\hat{\boldsymbol{w}}$ is sufficiently close to $V_{\infty}$, then sampling an unlabeled point $\boldsymbol{u}$ from $X$ whose corresponding cut is close to $\hat{\boldsymbol{w}}$ can have the effect of significantly reducing the size of the version space, hence accelerating the convergence. Following this intuition, we propose a sampling strategy alternative to SIM, named Closest cutting plane (CUT). Given $\hat{\boldsymbol{w}}$ in the closure $\bar{V}_{t}$ of the current version space, CUT finds an unlabeled point $\boldsymbol{u}\in X$ whose corresponding hyperplane $h_{\boldsymbol{u}}=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{w}\cdot\boldsymbol{u}=1\}$ has minimum Euclidean distance from $\hat{\boldsymbol{w}}$. To accomplish this task, we solve the following mixed-integer quadratic program:
@@
$\ \boldsymbol{u}\in X,$ (4e)
$\ \boldsymbol{p}\in\mathbb{R}^{d}.$ (4f)
The distance between $\hat{\boldsymbol{w}}$ and the hyperplane $h_{\boldsymbol{u}}$ is modeled as the distance between $\hat{\boldsymbol{w}}$ and its projection $\boldsymbol{p}$ on $h_{\boldsymbol{u}}$, which is minimized by the objective function 4a. The quadratic constraint (4b) ensures that $\boldsymbol{p}$ belongs to $h_{\boldsymbol{u}}$. No-good cuts (4c) and (4d) are used exactly as in SIM.
@@ 3.2 Approximation of a Separation Oracle and the Resulting Linear Separation Algorithm
In principle, given any vector of surrogate weights in $V_{\infty}$, one can solve the original combinatorial optimization problem (1) by replacing the unknown constraint with the surrogate constraint. Therefore, in order to solve (1), one can try to find a point in the convex set $V_{\infty}$ using a membership oracle of the constraint $\bar{w}\cdot\boldsymbol{x}\leq 1$ with respect to $X$. It is well-known that one can efficiently find a point in a well-bounded convex body $K$ with a separation oracle. Specifically, assuming that a compact convex set $K\subseteq\mathbb{R}^{d}$ is contained in the Euclidean ball of radius $R$ centered at the origin, there exist algorithms using $O(d\log(dR/\epsilon))$ calls to the separation oracle to either find a point in $K$ or prove that $K$ does not contain any Euclidean ball of radius $\epsilon$ [8, 20, 25]. However, such algorithms are not directly applicable in our case because we do not have direct access to a separation oracle of $V_{\infty}$. On the other hand, given a weak separation oracle of a convex body $K$ (that finds an almost violated linear inequality valid for $K$), there exist algorithms that can find a point almost in $K$ in oracle-polynomial time [15]. To hopefully find a point in $V_{\infty}$ (or at least get a good approximation of $V_{\infty}$), we propose to construct an approximate separation oracle of $V_{\infty}$ using the membership oracle of (1b), and replace the separation oracle in existing convex optimization algorithms by the constructed approximate separation oracle. In particular, we adapt the algorithm proposed in [8] as our choice of the convex optimization algorithm in the oracle model. Implementation details are described below.
@@
The original version of the algorithm presented in [8] is intended to minimize a convex function $f$ over a full-dimensional convex body $K\subseteq\mathbb{R}^{d}$. Specifically, it generates a sequence of points in $\mathbb{R}^{d}$ that converges to a minimum inside $K$, under the assumption that $K$ is contained inside a ball of radius $R$ centered in the origin. Such a set $K$ is given implicitly by means of a separation oracle, which for every point $\hat{\boldsymbol{w}}\in\mathbb{R}^{d}$, either states that $\hat{\boldsymbol{w}}\in K$ or returns an inequality $\boldsymbol{\alpha}\cdot\boldsymbol{w}\leq\beta$ valid for $K$ but violated by $\hat{\boldsymbol{w}}$. To address the solution of problem 1, we assume $f$ to be constant and we let $K$ be the set $V_{\infty}$ defined in the previous Section 3.1. Since we do not have access to a separation oracle for $K$, we approximate it using the membership oracle at our disposal, that characterizes the set $S^{+}_{\infty}$ of all the feasible solutions. For each feasible point $\boldsymbol{s}^{+}$, the inequality $\hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{+}\leq 1$ is valid for $K$. Similarly, for each infeasible point $\boldsymbol{s}^{-}$, the inequality $-\hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{-}\leq-1$ is also valid for $K$. For every pair of sets $S^{+}_{t}$ and $S^{-}_{t}$ of labeled points, the valid inequalities corresponding to such points in addition to the inequalities that define $W$ (i.e., $\boldsymbol{a}\cdot\boldsymbol{w}\leq b$ for $(\boldsymbol{a},b)\in\Delta$) determine the current version space $V_{t}\supseteq K$ as defined in Section 3.1.
[…]
@@
The original algorithm from [8] guarantees a convergence rate towards a vector in $K$. However, in our case, no guarantees can be given regarding the convergence of $\hat{\boldsymbol{w}}_{t}$ towards a vector in $K=V_{\infty}$, because we are approximating the separation oracle required by the original algorithm using a membership oracle. We base the evaluation of the adapted algorithm on the experimental results in Section 4.
4 Experimental Results
In this section, we evaluate different combinations of linear separator methods and sampling strategies in the application of Algorithm 1. In particular, we combine SVM with SIM as traditionally done in active learning, then we apply SVM with our sampling strategy CUT and finally we test our alternative linear separator SEP with CUT. We evaluate such three combinations on two test problems, namely the pure knapsack problem, in which $X=\{0,1\}^{d}$, and the college study plan problem. We show how differences arise between the combinations in terms of the objective value of the best found feasible solutions.
@@ 4.1 College Study Plan Problem
The college study plan problem (CSPP) is a 0-1 combinatorial optimization problem in which a prospective college student needs to design their study plan in order to maximize the number of credits acquired respecting an estimated time budget. Each course yields a predefined number of credits and requires an effort in terms of study hours given by a certain fixed number and an amount of hours for independent study that depends on the particular student. Given a possible study plan, the student needs to decide whether or not taking the courses specified in the plan is feasible with respect to the time budget. We assume that an unknown knapsack constraint models the satisfaction of the time budget. Additional known constraints model the dependencies between the available courses present in the course catalog. Any course $i$ can be included in a study plan only if at least one course from each of a sequence of prerequisite sets $P^{i}_{1}$, $P^{i}_{2}$, $\ldots$, $P^{i}_{N^{i}}$ is included in the plan. Every course $i$ also has a set $C^{i}$ of corequisites (courses that have to be taken if and only if course $i$ is taken) and a set $A^{i}$ of alternatives (courses that must be taken if $i$ is taken). Reasonable assumptions avoid inconsistency of the dependencies, namely: (a) $\bigcup_{k=1}^{N^{i}}P^{i}_{k}$, $C^{i}$ and $A^{i}$ are pairwise disjoint sets for every $i\in\{1,\ldots,d\}$; (b) if $i\in\bigcup_{k=1}^{N^{j}}P^{j}_{k}$ then $j\notin\bigcup_{k=1}^{N^{i}}P^{i}_{k}$, for every $i,j\in\{1,\ldots,d\}$; (c) $i\in C^{j}$ if and only if $j\in C^{i}$ for every $i,j\in\{1,\ldots,d\}$; (d) $i\in A^{j}$ if and only if $j\in A^{i}$ for every $i,j\in\{1,\ldots,d\}$. The goal is to solve the following 0-1 integer linear programming model with an unknown knapsack constraint:
@@ 4.2 Computational Results
The query limit was set equal to 2000 for all the experiments. Other experimental setups are presented in Appendix C. The algorithms were tested on 3 sets of instances. We briefly describe their structure and we refer to Appendix B for the generation details. First, we obtained a set of knapsack instances with $d=100$ items and 3 types of rewards and weights: uncorrelated (KNAP-U), weakly correlated (KNAP-W) and strongly correlated (KNAP-S). For each type, we produced 30 instances. For CSPP we generated two sets of instances of dimension $d=150$. The first one (MIT-Math) contains 30 instances structured according to the list of undergraduate courses in Mathematics at the Massachusetts Institute of Technology (MIT). The second one (SYNTH-MIT) is composed by additional 30 synthetic instances, with randomly generated structures inspired by the dependencies of the previous MIT courses.
For different sets of instances, we summarize in Figures 1-5 how the relative objective error, averaged over the specific instances, decreases as the number of used oracle calls increases. Here the relative objective error is computed by $(z^{*}-\boldsymbol{v}\cdot\hat{\boldsymbol{x}})/z^{*}\cdot 100\%$ with $\hat{\boldsymbol{x}}$ denoting the best feasible solution found by individual methods. We first note that the average performance of all test methods on randomly generated instances is significantly better than the worst case lower bound presented in Example 1. Even though the dimension of the test problems is at least 100, all test methods can efficiently find feasible solutions with average relative objective error lower than 10% using up to a few hundred oracle calls. In particular, the relative objective error usually drops quickly at the very beginning and then tails off. Unlike traditional knapsack problems [31], strongly correlated knapsack instances seem easier, in terms of the number of oracle queries needed, to find almost optimal solutions with small (e.g., 1%) relative objective error, than uncorrelated and weakly correlated knapsack instances. The overall performance of our newly proposed methods is very satisfactory. On the knapsack instances, the method SEP+CUT outperforms the other alternatives (and SVM+CUT outperforms SVM+SIM) in relative objective error when the 2000 query limit is reached. On the CSPP instances, SEP+CUT converges to 1% relative objective error much faster than the other alternatives even though SVM+SIM attains a smaller average ending relative objective gap on MIT-Math CSPP instances. Additional experimental results are presented in Appendix D.
@@ 5 Conclusions
We proposed an active learn-and-optimize framework and its alternative realizations, for solving combinatorial optimization problems with an unknown linear constraint accessed by a membership oracle. Extensive numerical experiments are conducted to test the effectiveness of our proposed approaches. Even though such problems can be very hard in the worst case, we find that “average" combinatorial instances are not much harder than its convex counterpart. An interesting open question is how to balance the frequency of sampling and optimization steps. It would also be interesting to see if there exists any algorithm that matches the lower bound result in Example 1.
Appendix A Hardness Example
Example 1.
For any fixed $\epsilon$ such that $1/\epsilon$ is an integer constant, consider a knapsack problem instance with $n$ items, where $1/\epsilon$ items have weights equal to $1$, $(n-1/\epsilon)$ items have weights equal to $1/(1-\epsilon)$, and all items have identical profit equal to $1$. Assume that the knapsack has capacity $1/\epsilon$, and the item weights are unknown such that the knapsack constraint can only be accessed through a membership oracle. This knapsack instance has a unique optimal solution by picking all the weight $1$ items, which attains the optimal objective value $1/\epsilon$. Note that all other feasible solutions pick at most $11/\epsilon-$ 1 items and all solutions that pick at most $11/\epsilon-$ 1 items are feasible. Therefore, one cannot distinguish the weight $1$ items from the weight $1/(1-\epsilon)$ items from the oracle calls until one calls the membership oracle to query the feasibility of the optimal solution that picks all $1/\epsilon$ weight 1 items, in the sense that any solution that picks at most $11/\epsilon-$ 1 item is feasible while any solution that picks at least $1/\epsilon$ items is infeasible except the optimal solution. To find any better-than-$(1-\epsilon)$ solution (i.e., to find the optimal solution), any algorithm would in expectation (with respect to a random permutation of the order of the items) have to query the feasibility of almost all solutions that pick exactly $1/\epsilon$ items, which means at least $\Omega\left({{n}\choose{1/\epsilon}}\right)=\Omega(n^{1/\epsilon})$ calls to the membership oracle of the knapsack constraint.
Appendix B Instance Generation

### v3

@@ 3.1 Closest Cutting Plane Sampling Strategy
For the pair of sets $S^{+}_{t}$ and $S^{-}_{t}$ of feasible and infeasible points at time step $t$, the current version space is defined as the set
@@
Let us also define $S^{+}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1\}$ and $S^{-}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}>1\}$, as the sets of all feasible and infeasible points, respectively. Every time we update $S^{+}_{t}$ and $S^{-}_{t}$ with new labeled points to obtain $S^{+}_{t+1}$ and $S^{-}_{t+1}$, the next version space $V_{t+1}$ will be defined by additional inequalities with respect to $V_{t}$. The sequence of version spaces converges (in a finite but potentially exponential number of steps) to the set of surrogate weights whose corresponding surrogate constraint classifies every point in $X$ on the same side as the unknown knapsack constraint $\bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$, namely,
@@
=$=\{\boldsymbol{w}\in W\ |\ {\boldsymbol{w}}\cdot\boldsymbol{s}^{+}\leq 1\ \text{for all}\ \boldsymbol{s}^{+}\in S^{+}_{\infty}\text{,}\ \boldsymbol{w}\cdot\boldsymbol{s}^{-}>1\ \text{for all}\ \boldsymbol{s}^{-}\in S^{-}_{\infty}\}\text{.}$ (3)
Intuitively, if the vector of surrogate constraints $\hat{\boldsymbol{w}}$ is sufficiently close to $V_{\infty}$, then sampling an unlabeled point $\boldsymbol{u}$ from $X$ whose corresponding cut $\boldsymbol{w}\cdot\boldsymbol{u}=1$ is close to $\hat{\boldsymbol{w}}$ can have the effect of significantly reducing the size of the version space, hence accelerating the convergence. Following this intuition, we propose a sampling strategy alternative to SIM, named closest cutting plane (CUT). Given $\hat{\boldsymbol{w}}$ in the closure $\bar{V}_{t}$ of the current version space, CUT finds an unlabeled point $\boldsymbol{u}\in X$ whose corresponding hyperplane $h_{\boldsymbol{u}}=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{w}\cdot\boldsymbol{u}=1\}$ has minimum Euclidean distance from $\hat{\boldsymbol{w}}$. To accomplish this task, we solve the following mixed-integer quadratic program:
@@
$\ \boldsymbol{u}\in X,~{}\boldsymbol{p}\in\mathbb{R}^{d}.$ (4e)
The distance between $\hat{\boldsymbol{w}}$ and the hyperplane $h_{\boldsymbol{u}}$ is modeled as the distance between $\hat{\boldsymbol{w}}$ and its projection $\boldsymbol{p}$ on $h_{\boldsymbol{u}}$, which is minimized by the objective function (4a). The quadratic constraint (4b) ensures that $\boldsymbol{p}$ belongs to $h_{\boldsymbol{u}}$. No-good cuts (4c) and (4d) are used exactly as in SIM.
@@ 3.2 Approximation of a Separation Oracle and the Resulting Linear Separation Algorithm
In principle, given any vector of surrogate weights in $V_{\infty}$, one can solve the original combinatorial optimization problem (1) by replacing the unknown constraint with the surrogate constraint. Therefore, in order to solve (1), one can try to find a point in the convex set $V_{\infty}$ using a membership oracle of the constraint $\bar{w}\cdot\boldsymbol{x}\leq 1$ with respect to $X$. It is well known that one can efficiently find a point in a well-bounded convex body $K$ with a separation oracle. Specifically, assuming that a compact convex set $K\subseteq\mathbb{R}^{d}$ is contained in the Euclidean ball of radius $R$ centered at the origin, there exist algorithms using $O(d\log(dR/\epsilon))$ calls to the separation oracle to either find a point in $K$ or prove that $K$ does not contain any Euclidean ball of radius $\epsilon$ [8, 20, 25]. However, such algorithms are not directly applicable in our case because we do not have direct access to a separation oracle of $V_{\infty}$. On the other hand, given a weak separation oracle of a convex body $K$ (that finds an almost violated linear inequality valid for $K$), there exist algorithms that can find a point almost in $K$ in oracle-polynomial time [15]. To hopefully find a point in $V_{\infty}$ (or at least get a good approximation of $V_{\infty}$), we propose to construct an approximate separation oracle of $V_{\infty}$ using the membership oracle of (1b), that characterizes the set $S^{+}_{\infty}$ of all the feasible solutions, and replace the separation oracle in existing convex optimization algorithms by the constructed approximate separation oracle. In particular, we adapt the algorithm proposed in [8] as our choice of the convex optimization algorithm in the oracle model.
@@
The original version of the algorithm presented in [8] is intended to minimize a convex function $f$ over a full-dimensional convex body $K\subseteq\mathbb{R}^{d}$. Specifically, it generates a sequence of points in $\mathbb{R}^{d}$ that converges to a minimum inside $K$, under the assumption that $K$ is contained inside a ball of radius $R$ centered in the origin. Such a set $K$ is given implicitly by means of a separation oracle, which for every point $\hat{\boldsymbol{w}}\in\mathbb{R}^{d}$, either states that $\hat{\boldsymbol{w}}\in K$ or returns an inequality $\boldsymbol{\alpha}\cdot\boldsymbol{w}\leq\beta$ valid for $K$ but violated by $\hat{\boldsymbol{w}}$. To address the solution of problem (1), we assume $f$ to be constant and we let $K$ be the set $V_{\infty}$ defined in Section 3.1. To approximate a separation oracle for $K$ by using the membership oracle at our disposal, we collect a valid inequality for $K$ every time we label a point regardless of whether the current surrogate weights $\hat{\boldsymbol{w}}$ violate the inequality or not. For each feasible point $\boldsymbol{s}^{+}$, the inequality is $\boldsymbol{w}\cdot\boldsymbol{s}^{+}\leq 1$. For each infeasible point $\boldsymbol{s}^{-}$, the inequality is $-\boldsymbol{w}\cdot\boldsymbol{s}^{-}\leq-1$. This approximation is designed to be effective when using CUT as a sampling strategy, as the intuition is that if we can obtain objective value 00 in (4), then the valid inequality corresponding to the sampled point separates $\hat{\boldsymbol{w}}$ from $K$.
As in [8], let us define the norm $\|(\boldsymbol{x},y)\|=\sqrt{2}\|(x_{1}/R,\ldots,x_{d}/R,1)\|_{2}$ and the potential $\Phi(\boldsymbol{\alpha},\beta)=\frac{1}{4}\|(R\boldsymbol{\alpha},\beta)\|_{2}^{2}$, with $R=\sqrt{d}$, the diameter of $[0,1]^{d}$. The pseudo-code for the resulting algorithm, that we call SEP, is given in Algorithm 2. Set $\Lambda$ collects all currently available valid inequalities for $K$, which are the same that describe the current version space $V_{t}\supseteq K$ as defined in Section 3.1 (including those that define $W$). Each inequality $\boldsymbol{\alpha}\cdot\boldsymbol{w}\leq\beta$ is encoded as a pair $(\boldsymbol{\alpha},\beta)$ and divided by the norm $\|(\boldsymbol{\alpha},\beta)\|$ of the pair.
@@
The original algorithm from [8] guarantees a convergence rate towards a vector in $K$. However, in our case, no guarantee can be given regarding the convergence of $\hat{\boldsymbol{w}}_{t}$ towards a vector in $K=V_{\infty}$, because we are approximating the separation oracle required by the original algorithm using a membership oracle. Therefore, we focus on the empirical performance of the algorithm rather than on theoretical guarantees.
4 Numerical Experiments
In this section, we evaluate different combinations of linear separator methods and sampling strategies in the application of Algorithm 1. In particular, we combine SVM with SIM as traditionally done in active learning, then we apply SVM with our sampling strategy CUT and finally we test our alternative linear separator SEP with CUT. We evaluate such three combinations on two test problems, namely the pure knapsack problem, in which $X=\{0,1\}^{d}$, and the college study plan problem, as a variant of [1]. We show how differences arise between the combinations in terms of the objective value of the best found feasible solutions.
@@ 4.1 College Study Plan Problem
We first briefly describe the second test problem, i.e., the college study plan problem, as a variant of the personalized learning problem in [1]. The college study plan problem (CSPP) is a 0-1 combinatorial optimization problem in which a prospective college student needs to design their study plan in order to maximize the number of credits acquired respecting an estimated time budget. Each course yields a predefined number of credits and requires an effort in terms of study hours given by a certain fixed number and an amount of hours for independent study that depends on the particular student. Given a possible study plan, the student needs to decide whether or not taking the courses specified in the plan is feasible with respect to the time budget. We assume that an unknown knapsack constraint models the satisfaction of the time budget. Additional known constraints model the dependencies between the available courses present in the course catalog. Any course $i$ can be included in a study plan only if at least one course from each of a sequence of prerequisite sets $P^{i}_{1}$, $P^{i}_{2}$, $\ldots$, $P^{i}_{N^{i}}$ is included in the plan. Every course $i$ also has a set $C^{i}$ of corequisites (courses that have to be taken if and only if course $i$ is taken) and a set $A^{i}$ of alternatives (courses that cannot be taken if $i$ is taken). The goal is to solve the following problem with unknown $\bar{\boldsymbol{w}}$:
@@ 4.2 Computational Results
The query limit was set equal to 2000 for all the experiments and we initialized $S^{+}=\{\boldsymbol{0}\}$ and $S^{-}=\{\boldsymbol{1}\}$. Other experimental setups are presented in Appendix 0.B. Instance generation details are presented in Appendix 0.C. First, we obtained a set of knapsack instances with $d=100$ items and 3 types of rewards and weights: uncorrelated (KNAP-U), weakly correlated (KNAP-W) and strongly correlated (KNAP-S). For each type, we produced 30 instances. For CSPP, we generated two sets of instances of dimension $d=150$. The first one (MIT-Math) contains 30 instances structured according to the list of undergraduate courses in Mathematics at MIT. The second one (SYNTH-MIT) is composed by additional 30 synthetic instances, with randomly generated structures inspired by the dependencies of the previous MIT courses.
For different sets of instances, we summarize in Figures 1-5 how the relative objective error, averaged over the specific instances, decreases as the number of used oracle calls increases. Here, the relative objective error is computed by $(z^{*}-\boldsymbol{v}\cdot\hat{\boldsymbol{x}})/z^{*}\cdot 100\%$ with $\hat{\boldsymbol{x}}$ denoting the best feasible solution found by individual methods. We first note that the average performance of all test methods on randomly generated instances is significantly better than the worst case lower bound stated in Theorem 1.1. Even though the dimension of the test problems is at least 100, all test methods can efficiently find feasible solutions with average relative objective error lower than 10% using up to a few hundred oracle calls. In particular, the relative objective error usually drops quickly at the very beginning and then tails off. Unlike traditional knapsack problems [31], strongly correlated knapsack instances seem easier, in terms of the number of oracle queries needed, to find almost optimal solutions with small (e.g., 1%) relative objective error, than uncorrelated and weakly correlated knapsack instances. The overall performance of our newly proposed methods is strong. On the knapsack instances, the method SEP+CUT outperforms the other alternatives (and SVM+CUT outperforms SVM+SIM) in relative objective error when the 2000 query limit is reached. On the CSPP instances, SEP+CUT converges to 1% relative objective error much faster than the other alternatives even though SVM+SIM attains a smaller average ending relative objective gap on MIT-Math CSPP instances. Additional experimental results are presented in Appendix 0.D.
@@ 5 Conclusions
We proposed an active learn-and-optimize framework and its alternative realizations, for solving combinatorial optimization problems with an unknown knapsack constraint accessed by a membership oracle. We show that such problems can be very hard in the worst case. On the other hand, we empirically find that, under our framework, optimal solutions for average combinatorial instances can be approximated using a small number of oracle calls. An interesting open question is if there exists any algorithm that matches the lower bound result of Theorem 1.1, or if such a result can be improved.
Appendix 0.A Proof of Theorem 1.1
Proof
For any fixed $\epsilon\in(0,1)$, consider the instance $\mathcal{I}$ of the problem defined by
$\boldsymbol{v}=\boldsymbol{1}$,   $\bar{\boldsymbol{w}}=\dfrac{\boldsymbol{1}}{C}$,
where $C=\lceil 1/\epsilon\rceil-1$. The feasible solutions are those with at most $C$ items equal to 1. In particular, all the solutions picking $C$ items equal to 1 are feasible and all those picking $C+1$ items equal to 1 are infeasible. Consider now the set $\mathcal{P}$ composed of instances of the problem that satisfy
$\boldsymbol{v}^{\prime}=\boldsymbol{1}$,   $\bar{w}_{i}^{\prime}=\begin{cases}\dfrac{1}{C+1}&\text{for a subset of }C+1\text{ items},\\ \dfrac{1}{C}&\text{for all other items.}\end{cases}$
Each instance in $\mathcal{P}$ has a unique optimal solution by picking all the items with weight $1/(C+1)$, which attains the optimal objective value $\lceil 1/\epsilon\rceil=C+1$. All other solutions picking at least $C+1$ items equal to 1 are infeasible and all the solutions picking at most $C$ items equal to 1 are feasible. For every $\mathcal{I}^{\prime}\in\mathcal{P}$, the feasible sets for instance $\mathcal{I}$ and $\mathcal{I}^{\prime}$ coincide except that the set for $\mathcal{I}^{\prime}$ contains an additional solution, which is also optimal.
Note that, for a deterministic algorithm, the solution to be labeled can only be determined by the previously labeled solutions. Since the solution algorithm is not aware of the actual item weights, it cannot distinguish between $\mathcal{I}$ and any instance $\mathcal{I}^{\prime}\in\mathcal{P}$ unless the optimal solution of $\mathcal{I}^{\prime}$ is labeled. In particular, this is true even if the algorithm has already labeled an optimal solution for $\mathcal{I}$. For instance $\mathcal{I}$, we assume the solution algorithm does not terminate before it labels all solutions picking $C+1$ items (since before that the algorithm cannot guarantee the optimality of any labeled solution). Then, given the sequence of solutions that the algorithm labels for $\mathcal{I}$, one can construct an instance ′′$\mathcal{I}^{\prime\prime}\in\mathcal{P}$ that runs in exponential time for the algorithm. Let $\tilde{x}$ be the lastly labeled solution that picks $C+1$ items. Define the instance ′′$\mathcal{I}^{\prime\prime}\in\mathcal{P}$ as follows:
$\boldsymbol{v}^{\prime\prime}=\boldsymbol{1}$,   $\bar{w}_{i}^{\prime\prime}=\begin{cases}\dfrac{1}{C+1}&\text{if }\tilde{x}_{i}=1,\\ \dfrac{1}{C}&\text{if }\tilde{x}_{i}=0.\end{cases}$
By definition, $\tilde{x}$ is the unique better-than-(1-$\epsilon$)-optimal solution of $\mathcal{I}^{\prime\prime}$, and is the last labeled solution among all the solutions that pick $C+1$ items. Labeling it can require the algorithm to at least label all the solutions picking $C+1=\lceil 1/\epsilon\rceil$ items. In this case, the algorithm requires at least (${{d}\choose{\lceil 1/\epsilon\rceil}}=\Omega\left(d^{1/\epsilon}\right)$ calls to the membership oracle. ∎
Appendix 0.B Experimental Setups
Appendix 0.C Instance Generation